# Revisiting Reliability in Large-Scale Machine Learning Research Clusters

### v1

@@ III Understanding Lay of the Land of Large-Scale ML Training Clusters
Preemptions and Failure Cascades. A second-order effect of job failures is their effect on other, lower priority and likely smaller jobs—resulting in cascades [42]. In practice, large jobs tend to be higher priority jobs and small jobs are the lowest priority. By virtue of being high priority, large jobs are scheduled quickly by preempting the low priority jobs. When a large, high priority job fails due to hardware instability, Slurm is configured to reschedule it, possibly preempting hundreds of jobs in the process. The worst-case version of this is a crash loop, where a single job is configured to requeue on failures (e.g., by using exception handling in the submission script). In the period we observe, we see a 1024 GPU job NODE_FAIL and subsequently requeue 35 times, causing a total of 548 preemptions (over 7k GPUs). Such situations should be avoided, as they cause excessive churn in the cluster, resulting in lost goodput.
Preemptions are a second-order effect when considering job failures. In our clusters, to help ensure even the lowest priority jobs are able to make progress, preemptions can only occur after two hours of runtime. Nevertheless, without precise checkpointing, some work will be lost when job preemption occurs. Critically, large jobs 1) are expected to lose significant work and 2) fail more frequently (Figure 7), resulting in quadratic goodput costs as a function of job size. To estimate the impact of various sources of goodput loss on the overall cluster goodput, including preemptions occuring due to a rescheduled failed job, we assume that all jobs checkpoint hourly, giving an average of half an hour of lost work. Using the Slurm logs, we determine which jobs 1 received a NODE_FAIL (cluster-related issue) or a FAILED status that we attributed a hardware issue, 2 were preempted (PREEMPTED status) because of an instigating NODE_FAIL or FAILED job, and estimate the lost goodput (the minimum value of the jobs runtime and 30 minutes, multiplied by the number GPUs allocated to the job). Figure 8 shows that, as expected, most lost goodput ($y$-axis) from failures and second-order preemptions (in terms of wasted runtime, ignoring possibly large impact on resource fragmentation) on RSC-1 is due to large jobs at the scale of 2-4 thousand GPUs ($x$-axis). On RSC-2, moderate-sized jobs make up a higher fraction of the goodput loss due to differences in job makeup (see Figure 6). Absolute goodput loss for RSC-2 is also an order of magnitude smaller than for RSC-1, a consequence of differences in job makeup and failure rate. While optimizing large jobs is clearly important, 16% of the total lost goodput resulting from hardware failures is due to second-order preemptions, which come from jobs of much smaller sizes. These results indicate that the cluster as a whole is impacted beyond the failures themselves.
@@
In this section, we provide 1 an expected value formulation of ETTR based on training job parameters, job priority and cluster resource availability as inputs, and 2 a design space exploration study using job-level data to estimate ETTR for the RSC-1 and RSC-2 clusters. For 1, our formulation allows us to model a particular job’s reliability properties by making assumptions about checkpoint frequency and restart overheads. Note that expected ETTR, $\mathbb{E}[\text{ETTR{}}]$, is most useful for longer training runs—by the law of large numbers, longer training runs will tend to have observed ETTR values closer to this expectation. Using our analytical formulation for expected ETTR helps us quickly estimate and understand the impact of optimizations, e.g., what is the impact of halving failure rate? For 2, we continue using the prior parameters to as a tool for exploring the relative importance of different contributors to job overhead—exploring what the necessary requirements would be for reasonable ETTR ($\sim 0.90$) on the next generation of very large GPU training runs (using $\mathcal{O}(10^{5})$ GPUs).
Approximating $\mathbb{E}[\text{ETTR{}}]$ analytically: First, define $Q$ as time the job was eligible to be scheduled but was waiting in the job queue, $R$ as productive runtime, and $U$ as unproductive runtime. The wallclock time, $W=Q+R+U$, is the total time since the job was first eligible to be scheduled until it completes. We consider the intervals between checkpoints as $\Delta t_{cp}$, the time it takes to perform initialization tasks (e.g., loading checkpoints, etc.) as $u_{0}$ and the expected queue time both after submission and after every interruption as $q$ (we assume queue times are drawn i.i.d. and are not systematically shorter after an interruption). The number of nodes the job consumes is $N_{\text{nodes}}$ and the cluster failure rate $r_{f}$ is the expected number of failures per node-day of runtime. The MTTF for the job is $(N_{\text{nodes}}r_{f})^{-1}$.
@@
$\mathbb{E}[\text{ETTR{}}]\approx\frac{1-N_{\text{nodes}}r_{f}\left(u_{0}+\frac{\Delta t_{cp}}{2}\right)}{1+N_{\text{nodes}}r_{f}\left(q+\frac{u_{0}}{R}\left(q+u_{0}+\frac{\Delta t}{2}\right)\right)}$ (1)
Which, for long-running, high priority jobs where queue time is small and $R\gg q+u_{0}+\frac{\Delta t_{cp}}{2}$ simplifies to
$\mathbb{E}[\text{ETTR{}}]\approx 1-N_{\text{nodes}}r_{f}\left(u_{0}+\frac{\Delta t_{cp}}{2}\right)$ (2)
@@
For the RSC clusters, $r_{f}\approx 5\times 10^{-3}$ failures per GPU node day of runtime, $t_{cp}\approx 1~{}\mathrm{hr}=0.04~{}\mathrm{days}$ and $u_{0}\approx 5-20~{}\mathrm{mins}$, and $(N_{\text{nodes}}r_{f})^{-1}\gtrsim 0.1~{}\mathrm{day}$. Comparing to a Monte Carlo approach for computing the various expectations involved, even for large, long-running hypothetical jobs (e.g., 8k GPUs), we find that the approximation above is accurate to within $\sim$5%.
@@
Comparing to actual job runs: We compare the expected value formulation of ETTR above with observations of actual job runs observed on both clusters. A job run is a collection of jobs (some may have different Job IDs) that are part of the same training task. We assume $\Delta t_{cp}$ is 60 minutes and $u_{0}$ is 5 minutes. We focus on longer job runs with at least 24 hours of total training time, and for jobs that run with the highest-priority to understand ETTR of our highest priority jobs. Note that in calculating job run ETTR, we do not consider health checks here; we assume every job in the job run is interrupted by an infra failure, whether it was caught or not, this means that our data estimate of ETTR should be an underestimate.
@@
Figure 9: Comparing expected ETTR from aggregate cluster and partition-level statistics with average estimated ETTR from actual job runs, assuming 60 minutes checkpoint interval for all jobs and a restart overhead of 5 minutes. Also shown are 90% CI around the mean value for job run statistics. CI: Confidence Interval.
@@
To obtain cluster-level failure rates $r_{f}$ needed to compute $\mathbb{E}[\text{ETTR{}}]$, we count failures as taking all jobs (not just job runs) that use more than 128 GPUs which are assigned a NODE_FAIL status plus the number of jobs with a status of FAILED for which we can attribute a critical health check firing in the last 10 minutes of the job (or 5 minutes after completion). We then divide the number of failures by the number of node-days of runtime (the sum of runtime in days multiplied by the number of allocated nodes). We find nominally that RSC-1 has an $r_{f}$ of 6.50 failures per thousand node days and RSC-2 has a significantly lower $r_{f}$ of 2.34 failures per thousand node days. This finding is also corroborated by looking at the rate at which GPUs are swapped in the cluster – we find RSC-1 GPUs are swapped at $\sim$ 3 times the rate compared to RSC-2; both the GPU swap rate and failure rate differences may be due to differing workloads that tax GPUs on RSC-1 more heavily.
@@
Analysis of ETTR Results. Figure 9 shows our findings. Except for smaller job runs ($<64$ GPUs), the $\mathbb{E}[\text{ETTR{}}]$ and average measured job run ETTR agree fairly well. On RSC-1, the largest job runs ($>1024$ GPUs) have systematically higher ETTR ($>0.9$) than predicted by $\mathbb{E}[\text{ETTR{}}]$. This is due to actual wait times for these larger job runs being shorter than average.
In a hypothetical scenario where the RSC-1 cluster became dedicated to a single 16,000 GPU training run occupying the entire cluster, the expected ETTR would be 0.7 for a 60 minute checkpoint interval. Moving to a 5 minute checkpoint interval would increase expected ETTR to 0.93, illustrating the value of frequent checkpointing to insulate against interruptions (assuming checkpoint writes are non-blocking).
Looking towards the future: for future job runs on other clusters that require $\mathcal{O}(10^{5})$ GPUs, the MTTF implied by an RSC-1-like failure rate (6.50 failures per thousand node days) is $\sim$15 minutes. This means that an hour checkpoint interval is untenable. Figure 10 shows the tradeoff between failure rate and checkpoint interval, showing that e.g., for an RSC-1-like failure rate, a checkpoint interval of $\sim$7 minutes is necessary to have an $\mathbb{E}[\text{ETTR{}}]=0.5$, which increases to $\sim$21 minutes if failure rates are closer to RSC-2. To reach ETTR of 0.9 at an RSC-2 failure rate, you would need $\sim$2 minute checkpointing and $\sim$2 minute restart overhead.
Figure 10: Checkpoint and failure rate requirements for 100k GPU-scale job runs.
@@
Observation 10: The RSC clusters are highly efficient (ETTR $>$ 0.9) for their largest and highest priority jobs. 1+ day 2048-4096 GPU job runs on RSC-1 show an average ETTR of over 0.9 at a one-hour assumed checkpoint interval, despite both clusters being congested and shared resources with 2-hour minimum time-to-preemption requirements for even the lowest priority jobs. To reach ETTR of 0.9 for the largest feasible training runs on RSC-1 (8,000 GPUs, taking half of RSC-1), checkpoint intervals need to be 30 minutes if queue times are 1 minute after every failure. To reach ETTR of 0.9 for a 100,000 GPU training runs on a hypothetical cluster with an RSC-2-like failure rate, checkpointing intervals and restart overhead need to be $\sim$2 minutes.
@@ IV-A Identifying and Repairing Lemon Nodes
Figure 11 illustrates the distribution of the signals based on a 28-day data snapshot for RSC-1, which we use to set the thresholds for detection criteria. The x-axis represents the number of occurrence for the signal per GPU node, normalized from 0 to 1. The y-axis represents the cumulative and normalized number of GPU nodes that experienced each signal. We found that user reported signal on excl_jobid_count did not have a strong correlation with node failures, yet a large number of nodes were excluded by at least one job. This motivates us to proactively detect lemon nodes instead leaving the burden of lemon detection to machine learning developers.
@@ IV-B Making Network Fabric Resilient via Adaptive Routing
The failure characterization analysis illustrates the significance of failures caused by the Infiniband link errors. Just as servers may fail in various stages, for transient or permanently degraded components, network links may undergo similar behavior due to changes in physical or electrical properties. Highly parallel model training at-scale will inevitably encounter faulty network links. These can be links with high error rates, flapping behavior that transitions between up and down states, permanently down links or high congestion in multi-tenant environments. All of these can result in degraded performance for communication across the fabric.
@@ V Key Lessons and Research Opportunities
Having presented our experience in operating and optimizing our clusters, we view several opportunities to improve reliability, manage rising infrastructure complexity, and co-design solutions with frameworks and algorithms.
@@
Training Reliability: We present how we run health checks and historic analysis to find transient failures, which inhibit a reliable compute substrate, as well as lemon nodes, whose removal can improve large job completion rates by over 30%. Looking forward, we see significant opportunities in further exposing reliability information to the scheduler as well as distributed algorithms [22], such that work is partitioned to maximize reliability or goodput. Additionally, we note potential improvement in the network fabric itself in being resilient by, for example, being able to reconfigure its topology to route around failures, similar in spirit to what we present on Adaptive Routing (§IV-B). We therefore envision future infrastructure systems that attempt to make unreliability less noticeable rather than attempting to remove it altogether. We believe rethinking entire systems may be necessary when future GPU systems, such as the NVIDIA GB200 [8], will change the unit of repair from a server to a rack, creating incentives to avoiding downtime by coping with failure.
@@
Closer to the application side, recall that the ETTR of a training job is a function of latency overheads associated with failures. An effective way to improve ETTR is to minimize restart latency cost, in addition to reducing the probability of restarts itself. As observed in the prior work [25], certain operations, such as NCCL initialization, can scale poorly with the number of GPU nodes. Looking forward, it is, therefore, important for future software systems to support fast and reliable routines—optimizing the latency cost of restarts. We expect optimizations, such as, replacing MPI-like collectives entirely and making preflight hardware tests more efficient, as key future avenues.
@@
Debugging Tools: Once many significant failures are eliminated via health checks, the remaining failures often manifest with proximal failures that do not immediately suggest a root cause. NCCL timeouts are one of the most common symptoms of a failure whose root cause could be network infrastructure, buggy model code, or other stuck components. Regularly checking hardware infrastructure health (§II-C) reduces the frequency of NCCL timeouts by finding faulty network or hardware issues proactively before they manifest as NCCL kernels becoming stuck, but identifying the root causes of the remaining timeouts leads to new health checks or bug fixes.
@@
We can improve the success rate of training runs by retroactively identifying the root cause of a NCCL timeout, by comparing logged data across different ranks participating in the collective. For instance, by logging which ranks started each collective, and the dependencies between collectives, we can find the first collective where some ranks started the collective but others did not, and further investigate the missing ranks. If all ranks entered but did not leave a collective, we can examine the network traffic within the collective to identify which component did not send or receive an expected message. Removing the culprit rank or network components from future runs will reduce the likelihood of those runs hitting the same issue. Better diagnosis and debugging tools are needed for efficient, yet reliable large-scale distributed training. One can imagine extending existing management tools, such as IPMI [23], to deliver machine debug information in an out-of-band network and closing the attribution gap.
@@
Programming Models: One confounding factor for diagnosing NCCL timeouts is the SPMD programming model as implemented in PyTorch. If different ranks accidentally issue collectives such as all-reduce in the wrong order, the job will deadlock, resulting in a NCCL timeout. Debugging a NCCL timeout, therefore, first starts with determining if the training script was buggy, adding a confounding factor to tracking down infrastructure instability. Dynamically detecting incorrect programs and raising exceptions rather than deadlocking would improve stability. Alternatively, one can aim to eliminate the possibility of mismatched collectives entirely. For instance, Pathways [14], introduces a single point where communication is scheduled ensuring each machine schedules communication consistently.

### v2

@@ III Understanding Lay of the Land of Large-Scale ML Training Clusters
Preemptions and Failure Cascades. A second-order effect of job failures is their effect on other, lower priority and likely smaller jobs—resulting in cascades [58]. In practice, large jobs tend to be higher priority jobs and small jobs are the lowest priority. By virtue of being high priority, large jobs are scheduled quickly by preempting the low priority jobs. When a large, high priority job fails due to hardware instability, Slurm is configured to reschedule it, possibly preempting hundreds of jobs in the process. The worst-case version of this is a crash loop [52], where a single job is configured to requeue on failures (e.g., by using exception handling in the submission script). In the period we observe, we see a 1024 GPU job NODE_FAIL and subsequently requeue 35 times, causing a total of 548 preemptions (over 7k GPUs). Such situations should be avoided, as they cause excessive churn in the cluster, resulting in lost goodput.
Preemptions are a second-order effect when considering job failures. In our clusters, to help ensure even the lowest priority jobs are able to make progress, preemptions can only occur after two hours of runtime. Nevertheless, without precise checkpointing, some work will be lost when job preemption occurs. Critically, large jobs 1) are expected to lose significant work and 2) fail more frequently (Figure 7), resulting in quadratic goodput costs as a function of job size. To estimate the impact of various sources of goodput loss on the overall cluster goodput, including preemptions occuring due to a rescheduled failed job, we assume that all jobs checkpoint hourly (we find this is a typical checkpoint interval for larger jobs on the RSC clusters), giving an average of half an hour of lost work. Using the Slurm logs, we determine which jobs 1 received a NODE_FAIL (cluster-related issue) or a FAILED status that we attributed a hardware issue, 2 were preempted (PREEMPTED status) because of an instigating NODE_FAIL or FAILED job, and estimate the lost goodput (the minimum value of the jobs runtime and 30 minutes, multiplied by the number GPUs allocated to the job). Figure 8 shows that, as expected, most lost goodput ($y$-axis) from failures and second-order preemptions (in terms of wasted runtime, ignoring a possibly large impact on resource fragmentation) on RSC-1 is due to large jobs at the scale of 2-4 thousand GPUs ($x$-axis). On RSC-2, moderate-sized jobs make up a higher fraction of the goodput loss due to differences in job makeup (see Figure 6). Absolute goodput loss for RSC-2 is also an order of magnitude smaller than for RSC-1, a consequence of differences in job makeup and failure rate. While optimizing large jobs is clearly important, 16% of the total lost goodput resulting from hardware failures on RSC-1 is due to second-order preemptions, which come from jobs of much smaller sizes. These results indicate that the cluster as a whole is impacted beyond the failures themselves.
@@
In this section, we provide 1 an expected value formulation of ETTR based on training job parameters, job priority, and cluster resource availability as inputs, and 2 a design space exploration using job-level data to estimate ETTR for the RSC-1 and RSC-2 clusters. For 1, our formulation allows us to model a particular job’s reliability properties by making assumptions about checkpoint frequency as well as checkpoint write and restart overhead. Note that expected ETTR, $\mathbb{E}[\text{ETTR{}}]$, is most useful for longer training runs—by the law of large numbers, longer training runs will tend to have observed ETTR values closer to this expectation. Using our analytical formulation for expected ETTR helps us quickly estimate and understand the impact of optimizations, e.g., what is the impact of halving failure rate? For 2, we continue using the prior parameters as a tool for exploring the relative importance of different contributors to job overhead—exploring what the necessary requirements would be for reasonable ETTR ($\sim$0.90) on the largest feasible RSC-1 training runs (up to $\sim$2/3 of the cluster, or $\sim$12,000 GPUs) under the typical paradigm used today: checkpointing progress to disk and restoring upon a restart, without fault tolerant solutions like spare idle compute.
Approximating $\mathbb{E}[\text{ETTR{}}]$ analytically: First, define $Q$ as time the job was eligible to be scheduled but was waiting in the job queue, $R$ as productive runtime, and $U$ as unproductive runtime. The wallclock time, $W=Q+R+U$, is the total time since the job was first eligible to be scheduled until it completes. We consider the intervals between checkpoints as $\Delta t_{cp}$ with each checkpoint write incurring a constant time penalty of $w_{cp}$, the time it takes to perform initialization tasks (e.g., loading checkpoints, etc.) as $u_{0}$, and the expected queue time both after submission and after every interruption as $q$ (we assume queue times are drawn i.i.d. and are not systematically shorter after an interruption). The number of nodes the job consumes is $N_{\text{nodes}}$ and the cluster failure rate $r_{f}$ is the expected number of failures per node-day of runtime. The MTTF for the job is $(N_{\text{nodes}}r_{f})^{-1}$.
@@
$\mathbb{E}[\text{ETTR{}}]\gtrsim\frac{1-N_{\text{nodes}}r_{f}\left(u_{0}+\frac{\Delta t_{cp}}{2}\right)}{1+\frac{u_{0}+q}{R}+\frac{w_{cp}}{\Delta t_{cp}}+N_{\text{nodes}}r_{f}q\left(1+\frac{w_{cp}}{\Delta t_{cp}}-\frac{\Delta t_{cp}}{2R}\right)}$ (1)
Which, for long-running, high priority jobs where queue time is much smaller than the MTTF and $R\gg q+u_{0}+\frac{\Delta t_{cp}}{2}$ simplifies to
$\mathbb{E}[\text{ETTR{}}]\approx\frac{1-N_{\text{nodes}}r_{f}\left(u_{0}+\frac{\Delta t_{cp}}{2}\right)}{1+\frac{w_{cp}}{\Delta t_{cp}}}$ (2)
While a full derivation of the optimal checkpoint interval using the equation for ETTR above is possible, a classic result by Daly and Young [23, 63, 17, 16] shows that, under some limiting assumptions, the optimal checkpointing interval is approximately
$\Delta t_{cp}^{*}=\sqrt{\frac{2w_{cp}}{N_{\mathrm{nodes}}r_{f}}}$ (3)
@@
For the RSC clusters, $r_{f}\approx 5\times 10^{-3}$ failures per GPU node-day of runtime, $w_{cp}\approx 5~{}\mathrm{mins}$, $u_{0}\approx 5-20~{}\mathrm{mins}$, and $(N_{\text{nodes}}r_{f})^{-1}\gtrsim 0.1~{}\mathrm{day}$. Comparing to a Monte Carlo approach for computing the various expectations involved, even for large, long-running hypothetical jobs (e.g., 8k GPUs), we find that the approximation above is accurate to within $\sim$5%.
@@
Comparing to actual job runs: We compare the expected value formulation of ETTR above with observations of actual job runs observed on both clusters. A job run is a collection of jobs (some may have different Job IDs) that are part of the same training task. We assume $\Delta t_{cp}$ is Daly-Young optimal, and that $u_{0}$ and $w_{cp}$ are both 5 minutes. We focus on longer job runs with at least 48 hours of total training time and jobs that run with the highest priority. Note that in calculating job run ETTR, we do not consider health checks; we assume every job in the job run that does not exit cleanly (with a COMPLETED state assigned) is interrupted by an infra failure, whether it was caught or not, and this means that our data estimate of ETTR should be an underestimate.
@@
Figure 9: Comparing expected ETTR ($\mathbb{E}[\text{ETTR{}}]$) from aggregate cluster and partition-level statistics on both wait time and failure rate with average estimated ETTR from actual job runs, assuming Daly-Young optimal checkpointing with 5 minute restart overhead and 5 minute checkpoint write overhead. Error bars shown are 90% CI around the mean value of job run ETTR. CI: Confidence Interval is shown on empirical data.
@@
To obtain cluster-level failure rates $r_{f}$ needed to compute $\mathbb{E}[\text{ETTR{}}]$, we count failures as taking all jobs (not just job runs) that use more than 128 GPUs that are assigned a NODE_FAIL status plus the number of jobs with a status of FAILED for which we can attribute a critical health check firing in the last 10 minutes of the job (or 5 minutes after completion). We then divide the number of failures by the number of node-days of runtime (the sum of runtime in days multiplied by the number of allocated nodes). We find nominally that RSC-1 has an $r_{f}$ of 6.50 failures per thousand node-days and RSC-2 has a significantly lower $r_{f}$ of 2.34 failures per thousand node-days. This finding is also corroborated by looking at the rate at which GPUs are swapped in the cluster—we find RSC-1 GPUs are swapped at $\sim$3 times the rate compared to RSC-2; both the GPU swap rate and failure rate differences may be due to differing workloads that tax GPUs on RSC-1 more heavily.
@@
Analysis of ETTR Results. Figure 9 shows our findings. Our predictions of $\mathbb{E}[\text{ETTR{}}]$ and average measured job run ETTR agree fairly well, with measured job run ETTR being generally smaller than predicted due to our conservative assumption that every state besides COMPLETED indicates an infra-related interruption. On RSC-1, the largest job runs ($>1024$ GPUs) have systematically higher ETTR than predicted by $\mathbb{E}[\text{ETTR{}}]$. This is due to actual wait times for these larger job runs being shorter than average, possibly due to Slurm scheduling configurations that prefer larger jobs.
Looking towards the future: The largest jobs on RSC-1 today consume roughly a quarter of the cluster (4096 GPUs). High-priority research efforts may feasibly consume up to 2/3 of the cluster for a short amount of time. Figure 10 shows projected ETTR as a function of both failure rate and checkpoint write overhead. To obtain a good ETTR ($\gtrsim 0.9$) for a job of this size, RSC-1 failure rate either needs to improve from 6.50 to $\sim$1 or checkpoint write overhead needs to be under a minute $\mathcal{O}(10\mathrm{s})$, achievable with asynchronous checkpoint writing strategies [61], though note that our analytical model was created with a classical checkpoint strategy in mind.
Figure 10: Checkpoint and failure rate requirements for large 12k GPU-scale job runs. Contours interpolate between poor ETTR (0.7, light green) to almost perfect ETTR (0.99, dark green) as a function of cluster failure rate and checkpoint write overhead. Checkpoint intervals smaller than 10s shown in red.
@@
Observation 10: The RSC clusters are highly efficient (ETTR 0.85-0.9) for their largest and highest priority jobs. 2+ day 2048-4096 GPU job runs on RSC-1 show an average ETTR of around 0.9 assuming checkpoint intervals are Daly-Young optimal and that both restart overhead and checkpoint write overhead is 5 minutes, despite both clusters being congested and shared resources with 2-hour minimum time-to-preemption requirements for even the lowest priority jobs. To reach ETTR of 0.9 for the largest feasible training runs on RSC-1 ($\sim$12,000 GPUs, taking two thirds of RSC-1), checkpoint write time overhead needs to be on the order of $\sim$10 seconds or failure rate needs to dramatically improve from 6.50 to $\sim$1.
@@ IV-A Identifying and Repairing Lemon Nodes
Figure 11 illustrates the distribution of the signals based on a 28-day data snapshot for RSC-1, which we use to set the thresholds for detection criteria. The x-axis represents the number of occurrence for the signal per GPU node, normalized from 0 to 1. The y-axis represents the cumulative and normalized number of GPU nodes that experienced each signal. We found that user reported signal on excl_jobid_count did not have a strong correlation with node failures, yet a large number of nodes were excluded by at least one job. This motivates us to proactively detect lemon nodes instead of leaving the burden of lemon detection to ML developers.
@@ IV-B Making Network Fabric Resilient via Adaptive Routing
The failure characterization analysis illustrates the significance of failures caused by the Infiniband link errors. Just as servers may fail in various stages, for transient or permanently degraded components, network links may undergo similar behavior due to changes in physical or electrical properties. Highly parallel model training at-scale will inevitably encounter faulty network links. These can be links with high error rates, flapping behavior that transitions between up and down states, permanently down links, or high congestion in multi-tenant environments. All of these can result in degraded performance for communication across the fabric.
@@ V Key Lessons and Research Opportunities
This section summarizes lessons learned and highlight key opportunities to improve reliability, manage infrastructure complexity, and co-design solutions.
Training Reliability: As suggested by Observations 1, 2, and 4, keeping nodes and therefore jobs healthy is a major challenge. By Observations 7, 8, 9, and 10, the impact of unhealthy infrastructure is felt asymmetrically across job sizes and priorities due to job-size dependent failure rates along with scheduling dynamics. We present how we run health checks and historic analysis to find transient failures, which inhibit a reliable compute substrate, as well as lemon nodes, whose removal can improve large job completion rates by over 30% (Observation 11). Our analysis does not focus on Silent Data Corruption (SDC) errors [25, 18]. A common technique to guard against SDCs is to monitor abrupt changes in gradients or activations during model convergence [33, 45]. While understanding SDCs is important future work, our own anecdotal experience with operating a research clusters is that such events rarely warrant user complaints.
@@
Looking forward, we see significant opportunities in further exposing reliability information to the scheduler as well as distributed algorithms [34, 65], such that work is partitioned to maximize reliability or goodput. Additionally, we note potential improvement in the network fabric itself in being resilient by, for example, being able to reconfigure its topology to route around failures, similar in spirit to what we present on Adaptive Routing (§IV-B, Observation 12). We therefore envision future infrastructure systems that attempt to make unreliability less noticeable rather than attempting to remove it altogether. We believe rethinking entire systems may be necessary when future GPU systems, such as the NVIDIA GB200 [8], will change the unit of repair from a server to a rack, creating incentives to avoiding downtime by coping with failure.
@@
Closer to the application side, recall that the ETTR of a training job is a function of latency overheads associated with failures. An effective way to improve ETTR is to minimize checkpoint [61] and restart latency costs, while also reducing the probability of restarts due to failures. As observed in prior work [38], certain operations, such as NCCL initialization, can scale poorly with the number of GPU nodes. Looking forward, it is, therefore, important for future software systems to support fast and reliable routines—optimizing the latency cost of restarts. We expect optimizations, such as, replacing MPI-like collectives entirely and making preflight hardware tests more efficient, as key future avenues.
@@
Debugging Tools: Once many significant failures are eliminated via health checks, the remaining failures often manifest with proximal failures that do not immediately suggest a root cause, as suggested by Observations 3, 5, and 6. NCCL timeouts are one of the most common symptoms of a failure whose root cause could be network infrastructure, buggy model code, or other stuck components. Regularly checking hardware infrastructure health (§II-C) reduces the frequency of NCCL timeouts by finding faulty network or hardware issues proactively before they manifest as NCCL kernels becoming stuck. Identifying the remaining root causes of NCCL timeouts may require new health checks and tools.
@@
We can improve the success rate of training runs by retroactively identifying the root cause of a NCCL timeout, by comparing logged data across different ranks participating in the collective. By logging which ranks started each collective and the dependencies between collectives, we can find the first collective where some ranks started the collective but others did not, and we can further investigate the missing ranks. If all ranks entered but did not leave a collective, we can examine the network traffic within the collective to identify which component did not send or receive an expected message [11]. Removing the culprit rank or network components from future runs will reduce the likelihood of those runs hitting the same issue. Better diagnosis and debugging tools are needed for efficient, yet reliable large-scale distributed training. One can imagine extending existing management tools, such as IPMI [36], to deliver machine debug information in an out-of-band network and closing the attribution gap.
New Programming Models and Algorithms: This work primarily focuses on traditional SPMD programming models with traditional Bulk-Synchronous Parallel [57] semantics. Going forward, relaxing both of these constraints may unlock additional reliability, since such future programs may be able to more gracefully tolerate failures.
@@
On the system side, one confounding factor for diagnosing NCCL timeouts is the SPMD programming model as implemented in PyTorch. If different ranks accidentally issue collectives such as All-Reduce in the wrong order, the job will deadlock, resulting in a NCCL timeout. Debugging a NCCL timeout, therefore, first starts with determining if the training script was buggy, adding a confounding factor to tracking down infrastructure instability. Dynamically detecting incorrect programs and raising exceptions rather than deadlocking would improve stability. Alternatively, one can aim to eliminate the possibility of mismatched collectives entirely. For instance, Pathways [15], introduces a single point where communication is scheduled ensuring each machine schedules communication consistently. On the algorithmic side, there is renewed interest in algorithms that are asynchronous. Past systems embraced asynchronous gradient exchange as a performance optimization [24, 19, 51]. Current research trends look at this as a potentially cross-cluster federated learning problem [26].